OpenStudy (anonymous):

On the fast speed way some drivers reach a average speed of 2 miles per minute. write a formula that give the number of miles m that such a driver would travel in x minutes. how far would this driver travel in 35 minutes? the formula is m the distance is

OpenStudy (binary3i):

average speed= m/x = 2 so m=2x

OpenStudy (nurali):

m=2x where x is time in minutes 35*2=70 miles

OpenStudy (anonymous):

the distance is

OpenStudy (binary3i):

m is the distance where m= 2x and x is the time so two times 35 giving 70 miles